OpenStudy (anonymous):

Solve x divided by nine is less than or equal to negative three for x can you tell me each step i would do to solve the question

OpenStudy (anonymous):

do you mean x/9 <=-3x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x/9 \(\le \) -3x multiply both side by 9, then become x \(\le \) -27x add both side by 27x, then become x + 27x \(\le\) 0 28 x \(\le\) 0 finally, divided both side by 28, then become x \(\le\) 0

OpenStudy (anonymous):

do u understand @yoyogators ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

ur welcome..